A corporate bod has a 6% coupon that is payable semi-annually. The bond is selling in the secondary market for $910. Its par value is $1,000. What is its current yield?

2. What is your return if you invest $2,094now and then you sell the investment for $6,884 tow years from now? Assume the nominal risk free rate is 5%.
Answer #1

1). Current Yield = Annual coupon payment / Current Bond Price = [6% * $1,000] / $910 = $60 / $910 = 0.0659, or 6.59%

2). Return = [Value in 2 two years / Value now]1/n - 1

= [$6,884 / $2,094]1/2 - 1 = [3.2875]1/2 - 1 = 1.8131 - 1 = 0.8131, or 81.31%